# PLAME: Leveraging Pretrained Language Models to Generate Enhanced Protein Multiple Sequence Alignments

### v1

@@ Abstract
Protein structure prediction is essential for drug discovery and understanding biological functions. While recent advancements like AlphaFold have achieved remarkable accuracy, most folding models rely heavily on multiple sequence alignments (MSAs) to boost prediction performance. This dependency limits their effectiveness on low-homology proteins and orphan proteins, where MSA information is sparse or unavailable. To address this limitation, we propose PLAME, a novel MSA design model that leverages evolutionary embeddings from pretrained protein language models. Unlike existing methods, PLAME introduces pretrained representations to enhance evolutionary information and employs a conservation-diversity loss to enhance generation quality. Additionally, we propose a novel MSA selection method to effectively screen high-quality MSAs and improve folding performance. We also propose a sequence quality assessment metric that provides an orthogonal perspective to evaluate MSA quality. On the AlphaFold2 benchmark of low-homology and orphan proteins, PLAME achieves state-of-the-art performance in folding enhancement and sequence quality assessment, with consistent improvements demonstrated on AlphaFold3. Ablation studies validate the effectiveness of the MSA selection method, while extensive case studies on various protein types provide insights into the relationship between AlphaFold’s prediction quality and MSA characteristics. Furthermore, we demonstrate that PLAME can serve as an adapter achieving AlphaFold2-level accuracy with the ESMFold’s inference speed.
@@ 1 Introduction
Understanding complex and highly variable protein structures is a cornerstone of drug design, enabling breakthroughs in target identification, validation, and drug-target interaction studies baker2001protein ; khoury2014protein ; jumper2021alphafold . Recent advancements, such as AlphaFold, have revolutionized structural biology by achieving near-experimental accuracy for most protein structures and their complexes jumper2021alphafold ; ahdritz2024openfold ; abramson2024accurate . However, many of these state-of-the-art models rely heavily on evolutionary information encoded in Multiple Sequence Alignments (MSAs) as input jumper2021alphafold ; abramson2024accurate , making their predictive performance highly dependent on the quality of the MSAs. This dependency poses significant challenges for low-homology protein families and orphan proteins, where evolutionary information is sparse or unavailable kwon2021assessment ; webb2016comparative .
Traditional approaches enhancing low-homology and orphan protein structure prediction include physics-based and template-based methods. Physics-based methods simulate energy functions but are computationally expensive and limited by approximations rohl2004rosetta ; cornell1995amber . Template-based methods rely on sequence homology detection hildebrand2009fast ; finn2011hmmer , making them unsuitable for orphan proteins where evolutionary information is scarce. These limitations have motivated the development of data-driven approaches to improve MSA generation.
Figure 1: Taxonomy of existing MSA designers. Previous methods capture discrete-space distribution of MSAs, including sequence inpainting and prompt-based generation. PLAME serves as an evolutionary-space-inspired de novo generation method independent of sequence prompt.
@@
Recent MSA generation methods can be categorized into sequence space inpainting and prompt-based conditional generation approaches (Figure 1). Sequence inpainting methods like MSA Generator and EvoGen operate within sequence space by completing partial MSAs based on patterns from existing data zhang2023enhancing ; zhang2022few , aiming to reconstruct evolutionary constraints. Prompt-based conditional generation methods, such as MSAGPT and EvoDiff, utilize pretrained models conditioned on MSA prompts to generate new sequences chen2024msagpt ; alamdari2023protein . These AI-based approaches have shown potential in augmenting low-homology MSAs and enhancing structure prediction. As an alternative approach, models like ESMFold construct implicit evolutionary representations directly from single sequences using pretrained protein language models lin2023evolutionary . While this strategy circumvents the need for MSAs, these methods cannot utilize structural templates or enhanced homology information, leading to limited folding accuracy.
Despite recent advancements, MSA generation methods targeting the enhancement of protein folding still face two major challenges. First, these models heavily rely on existing MSA databases, which are biased toward well-represented protein families and provide limited data for low-homology or orphan proteins. This reliance hinders their ability to construct evolutionary spaces that capture the unique constraints of such proteins, reducing their effectiveness in structure prediction. Second, there is a significant gap in understanding how MSA characteristics influence folding model performance. Current methods lack efficient evaluation metrics directly tied to folding performance, creating a disconnect between MSA generation objectives and structural accuracy. Additionally, the response of different protein types to MSA quality varies considerably, yet remains insufficiently studied. While approaches like MSAGPT have attempted to integrate structure prediction feedback through fine-tuning chen2024msagpt , their high computational cost limits scalability. A more comprehensive understanding of the MSA-folding relationship would enable more targeted optimization strategies.
To address these limitations, we propose PLAME, a PLm Aligner for MSA Enhancement designed for low-homology protein structure enhancement. Our key contributions are as follows:
@@
  • Enriched evolutionary spaces: PLAME leverages pretrained protein language models to incorporate evolutionary embeddings as input, providing richer evolutionary information than sparse MSAs for low-homology and orphan proteins wang2024maape ; mcwhite2023leveraging ; hong2024fast . Furthermore, we introduce a novel conservation-diversity loss, which increases model sensitivity to conserved regions while maintaining sequence diversity, enabling the generation of biologically plausible MSAs optimized for high-quality alignments.
  • Comprehensive MSA quality exploration: We propose a novel sequence quality assessment metric that provides a new perspective for evaluating MSA quality and its relationship to folding performance. Based on these insights, we develop an MSA selection method called High-Fidelity Appropriate Diversity (HiFiAD) that improves folding accuracy by effectively screening MSAs that balance sequence fidelity and diversity. This approach offers both analytical understanding and practical improvement for protein structure prediction.
  • Extensive experimental validation: Our experimental results demonstrate the substantial enhancement on challenging low-homology and orphan protein datasets when evaluated with both AlphaFold2 and AlphaFold3 folding frameworks. Ablation studies confirm the effectiveness of the conservation-diversity loss and HiFiAD selection method in improving both MSA quality and folding performance. Additionally, we conduct comprehensive experiments on multiple protein systems, including general and de novo proteins, providing deeper insights into the relationship between MSA characteristics and folding performance.
@@ 3.2 Model Architecture
[…]
@@
[…]
@@ 3.3 Conservation-Diversity Training Loss
The position-specific weights $w_{l}$ are derived from the Position-Specific Scoring Matrix (PSSM) and reflect the conservation level at each position. These weights are normalized to the range $[1-\delta,1+\delta]$, where $\delta$ controls sensitivity to conservation. Specifically,
@@
During model training, we apply $\delta=0.5$, assigning higher weights to conserved positions and lower weights to less conserved ones. For a batch of $N$ sequences, the PCE loss averages over all sequences and positions:
@@ 3.4 MSA Selection Method – HiFiAD
where $B$ is the BLOSUM matrix (BLOSUM62 in our case). For proteins with existing MSAs (termed few-shot cases), we limit the total number of concatenated MSAs to reduce noise in the original MSA distribution. Specifically, we cap the number of MSA sequences $N_{\text{aug}_{\text{max}}}$ at $\max(16,2N)$. If fewer than $N_{\text{aug}_{\text{max}}}$ virtual MSAs are available, we retain all generated MSAs to ensure sufficient diversity for downstream analysis.
@@ 4 Experiment
For the training dataset, we use the PDB and UniClust30 subsets from the open-source OpenProteinSet as our data source ahdritz2024openproteinset . The pre-searched MSAs from OpenFold training are also included. We only retain data with at least 64 MSA sequences. To avoid overlap with the test cases, we removed sequences with over 90% similarity using MMSeqs based on UniClust30 clustering results mirdita2017uniclust ; steinegger2017mmseqs2 . This process yields an initial dataset of 293,979 samples, which are split into training and validation sets with a 90:10 ratio. For the test dataset, we adopt the curated test cases from MSAGPT chen2024msagpt , which consist of 200 protein samples from three benchmarks: CASP14&15, CAMEO haas2018continuous , and PDB berman2000protein . Any $>90\%$ redundancy between the test cases and training dataset is eliminated.
@@ 4.1 PLAME as High-quality MSA Designer
We conducted experiments across three AF2 folding modes and evaluated six folding-related metrics. Results are shown in Table 1. We analyze the quality of MSAs generated by different methods and their performance in various scenarios. 1) Performance Superiority: PLAME achieved state-of-the-art performance on most metrics across all modes. This demonstrates that PLAME-generated MSAs effectively enhance AF2 folding, particularly in scenarios involving low-homology or orphan proteins, where significant improvements in prediction accuracy were observed. In contrast, EvoDiff and MSAGPT struggled to produce high-quality MSAs under these conditions. By leveraging the evolutionary latent space provided by ESM-2, PLAME generated biologically meaningful virtual MSAs, outperforming baselines on key metrics such as pLDDT and RMSD. 2) Diminishing Returns: As more advanced folding configurations were employed (e.g., with more powerful AF2 models or the inclusion of templates), the benefits of MSA augmentation gradually decreased. In Mode3, where templates were used, certain metrics exhibit performance decrease. This is likely because template information already captures much of the necessary sequence and evolutionary information, reducing the marginal utility of additional MSA augmentation. Furthermore, the noise introduced by low-quality MSAs can be amplified in stronger baselines, further impacting performance. 3) Metric Discrepancies: The enhancement effects across different metrics were not entirely consistent. For instance, the trends for pLDDT and RMSD diverged. While pLDDT, as a confidence metric, reflects the model’s confidence in its predictions, RMSD measures the global deviation of the predicted structure from the ground truth. MSA augmentation improved local prediction quality (leading to higher pLDDT), but had limited impact on reducing global structural deviation (resulting in minimal RMSD improvement). Nevertheless, pLDDT alone is sufficient to demonstrate the high quality of the generated MSAs. 4) Noise Issues: EvoDiff and MSAGPT produced MSAs that consistently underperformed compared to the original AF2 MSAs across all modes. This was particularly evident in low-homology protein scenarios, where the generated MSAs failed to accurately capture evolutionary distributions and often included irrelevant sequences. When concatenated with the original AF2 MSAs, these low-quality sequences introduced additional noise, disrupting the model’s inference. This aligns with the hypothesis presented in Section 3.4, where low-quality MSAs were predicted to introduce noise, ultimately impairing folding performance. 5) Few/Zero-shot Consistency: In Few-shot scenarios, EvoDiff and MSAGPT performed relatively better due to the presence of initial MSAs that provided additional homology information to guide the generation process. In contrast, PLAME does not rely on initial MSA searches and directly generates high-quality MSAs through its evolutionary latent space. This enables PLAME to excel in Zero-shot scenarios while maintaining consistent performance in Few-shot cases. This advantage significantly reduces runtime costs while ensuring robust enhancements across all scenarios.
@@ Appendix A Proof of Theorem
[…]
@@
Re-expressing each column term as $-\sum_{a}P\log Q=H\!\bigl{(}P(l,\!\cdot)\bigr{)}+\operatorname{\text{KL}}\!\bigl{(}P(l,\!\cdot)\|Q_{\theta}(l,\!\cdot)\bigr{)}$, we obtain
@@
Based on the fact that $\theta^{\star}_{\mathrm{PCE}}$ minimizes $D_{w}$, denote $\Delta_{w}:=D_{w}(\theta^{\star}_{\mathrm{CE}})-D_{w}(\theta^{\star}_{\mathrm{PCE}})\geq 0$. By applying ()($\ast$)( ∗ ) to both optimal parameters and subtracting, we obtain:
@@
Since $\Delta_{w}\geq 0$ and $01-\delta>$ 0; it is strictly positive whenever $\Delta_{w}>0$, Therefore,
@@
Since $10<\tau<$ 1, this transformation always increases entropy unless $P$ is already uniform:
@@ Appendix B Training and Sampling Details
Appendix C Extensive Experimental Results
C.1 Case Study on Successful Designs
To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structrure visualization in Appendix F. We select representative cases collected from different datasets and range from different lengths to comprehensive evaluate the samples.

### v2

# PLAME: Lightweight MSA Design Advances Protein Folding From Evolutionary Embeddings

@@ Abstract
Protein structure prediction often hinges on multiple sequence alignments (MSAs), which underperform on low-homology and orphan proteins. We introduce PLAME, a lightweight MSA design framework that leverages evolutionary embeddings from pretrained protein language models to generate MSAs that better support downstream folding. PLAME couples these embeddings with a conservation–diversity loss that balances agreement on conserved positions with coverage of plausible sequence variation. Beyond generation, we develop (i) an MSA selection strategy to filter high-quality candidates and (ii) a sequence-quality metric that is complementary to depth-based measures and predictive of folding gains. On AlphaFold2 low-homology/orphan benchmarks, PLAME delivers state-of-the-art improvements in structure accuracy (e.g., lDDT/TM-score), with consistent gains when paired with AlphaFold3. Ablations isolate the benefits of the selection strategy, and case studies elucidate how MSA characteristics shape AlphaFold confidence and error modes. Finally, we show PLAME functions as a lightweight adapter, enabling ESMFold to approach AlphaFold2-level accuracy while retaining ESMFold-like inference speed. PLAME thus provides a practical path to high-quality folding for proteins lacking strong evolutionary neighbors.
@@ 1 Introduction
Understanding complex and variable protein structures is central to target identification, validation, and drug–target interaction studies in drug design baker2001protein ; khoury2014protein ; jumper2021alphafold . Recent advances such as AlphaFold have transformed structural biology, achieving near-experimental accuracy across a wide range of proteins and complexes jumper2021alphafold ; ahdritz2024openfold ; abramson2024accurate . However, most state-of-the-art folding pipelines depend critically on evolutionary information encoded in Multiple Sequence Alignments (MSAs) jumper2021alphafold ; abramson2024accurate . As a consequence, their accuracy tracks the quality and depth of the available MSAs, leading to failure modes for low-homology families and orphan proteins where evolutionary neighbors are sparse or absent kwon2021assessment ; webb2016comparative . In practice, while MSA depth (e.g., effective sequence count) often correlates with predicted confidence (such as pLDDT), this relationship becomes highly unstable in sparse regimes where a few noisy or misaligned sequences can dominate the signal.
Historically, two classes of techniques have been used to mitigate weak homology. Physics-based modeling seeks low-energy conformations under hand-crafted or learned force fields, but is often computationally intensive and limited by approximations in the energy landscape rohl2004rosetta ; cornell1995amber . Template-based approaches leverage homology detection and profile–profile alignment to transfer structural priors from known folds hildebrand2009fast ; finn2011hmmer , yet degrade in regimes with poor evolutionary signal and are thus ill-suited to orphan proteins. These limitations have motivated data-driven strategies focused on improving the MSA itself rather than only the downstream folding network.
@@
Recent MSA design methods fall broadly into two lines (Figure 1). Sequence-space inpainting approaches (e.g., MSA Generator, EvoGen) complete or augment partial alignments by learning patterns directly in the discrete sequence space, aiming to reconstruct evolutionary constraints from observed MSAs zhang2023enhancing ; zhang2022few . Prompt-based conditional generation (e.g., MSAGPT, EvoDiff) uses pretrained models conditioned on MSA-style prompts to synthesize additional sequences chen2024msagpt ; alamdari2023protein . These approaches can deepen alignments and improve folding when homologs exist. An orthogonal line of work bypasses explicit MSAs by constructing implicit evolutionary representations from single sequences via large protein language models, as in ESMFold lin2023evolutionary . While MSA-free models avoid the homology bottleneck, they also forgo explicit template usage and enhanced homology signals, which can cap ultimate folding accuracy in challenging regimes.
Figure 1: Taxonomy of MSA designers. Prior work models the discrete MSA distribution via sequence inpainting or prompt-based generation. PLAME operates in an evolutionary embedding space to de novo generate MSAs without requiring sequence prompts.
Despite progress, two gaps remain for MSA design aimed at improving folding. (i) Biased supervision: Methods trained primarily on extant MSA databases inherit coverage biases toward well-studied families, offering limited guidance for low-homology and orphan proteins. This restricts the learned evolutionary manifold and hampers generalization exactly where better MSAs are most needed. (ii) Weak alignment–folding linkage: There is limited understanding—and few lightweight metrics—connecting MSA characteristics to downstream folding outcomes. As a result, generation objectives (e.g., likelihood in sequence space) can be misaligned with what most improves structure accuracy. While fine-tuning folding models with generated data (e.g., MSAGPT) partially closes the loop chen2024msagpt , the compute cost is high and does not yield general, model-agnostic criteria for selecting helpful MSAs. Closing this loop by retraining or adapting a folding model for each candidate MSA is prohibitive at scale, preventing iterative design-and-select cycles across large target sets.
We introduce PLAME (PLm Aligner for MSA Enhancement), a lightweight MSA designer targeted at low-homology and orphan proteins. PLAME departs from discrete sequence modeling and instead generates in an evolutionary embedding space derived from pretrained protein language models, then decodes to MSAs for downstream folding. This design enables PLAME to synthesize plausible evolutionary neighborhoods even when observed homologs are scarce, while remaining compatible with template-using pipelines.
@@
  • Enriched evolutionary spaces. PLAME conditions on evolutionary embeddings from pretrained protein language models wang2024maape ; mcwhite2023leveraging ; hong2024fast , providing richer contextual signals than sparse MSAs for low-homology targets. We introduce a conservation–diversity loss that explicitly balances agreement at conserved positions with appropriate sequence variability, steering generation toward biologically plausible and alignment-friendly MSAs.
  • Linking alignment to folding. We propose a complementary sequence-quality assessment metric that captures alignment properties predictive of folding gains, and we operationalize these insights with HiFiAD (High-Fidelity Appropriate Diversity), an MSA selection strategy that filters candidates to maximize downstream structure accuracy.
  • Broad, model-agnostic validation. Across challenging low-homology and orphan datasets, PLAME improves folding accuracy with both AlphaFold2 and AlphaFold3. Ablations isolate the effects of the conservation–diversity loss and HiFiAD on MSA quality and folding outcomes, and case studies spanning general and de novo proteins elucidate how specific MSA characteristics modulate AlphaFold confidence and error modes.
@@ 3.2 Model Architecture
$\text{Row-Attn}(\mathbf{Q}_{r},\mathbf{K}_{r},\mathbf{V}_{r})=\text{softmax}\left(\frac{\mathbf{Q}_{r}\mathbf{K}_{r}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{r}.$ (5)
@@
$\text{Col-Att}(\mathbf{Q}_{\text{c}},\mathbf{K}_{\text{c}},\mathbf{V}_{\text{c}})=\left(\text{softmax}\left(\frac{\mathbf{Q}_{\text{c}}\mathbf{K}_{\text{c}}^{\top}}{\sqrt{h}}\right)\mathbf{V}_{\text{c}}\right)^{\top}$ (7)
@@ 3.3 Conservation-Diversity Training Loss
The position-specific weights $w_{l}$ are derived from the Position-Specific Scoring Matrix (PSSM) and reflect the conservation level at each position. These weights are normalized to the range $1-\delta,1+\delta]$, where $\delta$ controls sensitivity to conservation. Specifically,
@@
where freq denotes the residue-frequency of 20 types of amino acids. During model training, we apply $\delta=0.5$, assigning higher weights to conserved positions and lower weights to less conserved ones. For a batch of $N$ sequences, the PCE loss averages over all sequences and positions:
@@ 3.4 MSA Selection Method – HiFiAD
where $B$ is the BLOSUM matrix (BLOSUM62 in our case); $s_{j}$ represents the j-th amino acid of sequence S with length L, and $m_{ij}$ refers to the j-th position of the i-th MSA sequence with length L. For proteins with existing MSAs (termed few-shot cases), we limit the total number of concatenated MSAs to reduce noise in the original MSA distribution. Specifically, we cap the number of MSA sequences $N_{\text{aug}_{\text{max}}}$ at $\max(16,2N)$. If fewer than $N_{\text{aug}_{\text{max}}}$ virtual MSAs are available, we retain all generated MSAs to ensure sufficient diversity for downstream analysis.
@@ 4 Experiment
For the training dataset, we use the PDB and UniClust30 subsets from the open-source OpenProteinSet as our data source ahdritz2024openproteinset . The pre-searched MSAs from OpenFold training are also included. We only retain data with at least 64 MSA sequences. To avoid overlap with the test cases, we removed sequences with over 90% similarity using MMSeqs based on UniClust30 clustering results mirdita2017uniclust ; steinegger2017mmseqs2 . This process yields an initial dataset of 293,979 samples, which are split into training and validation sets with a 90:10 ratio. For the test dataset, we adopt the curated test cases from MSAGPT chen2024msagpt , which consist of 200 protein samples from three benchmarks: CASP14&15, CAMEO haas2018continuous , and PDB berman2000protein . Any $90\%$ redundancy between the test cases and training dataset is eliminated.
@@ 4.1 PLAME as High-quality MSA Designer
We conducted experiments across three AF2 folding modes and evaluated six folding-related metrics. Results are shown in Table 3.4. We analyze the quality of MSAs generated by different methods and their performance in various scenarios. 1) Performance Superiority: PLAME achieved state-of-the-art performance on most metrics across all modes. This demonstrates that PLAME-generated MSAs effectively enhance AF2 folding, particularly in scenarios involving low-homology or orphan proteins, where significant improvements in prediction accuracy were observed. In contrast, EvoDiff and MSAGPT struggled to produce high-quality MSAs under these conditions. By leveraging the evolutionary latent space provided by ESM-2, PLAME generated biologically meaningful virtual MSAs, outperforming baselines on key metrics such as pLDDT and RMSD. 2) Diminishing Returns: As more advanced folding configurations were employed (e.g., with more powerful AF2 models or the inclusion of templates), the benefits of MSA augmentation gradually decreased. In Mode3, where templates were used, certain metrics exhibit performance decrease. This is likely because template information already captures much of the necessary sequence and evolutionary information, reducing the marginal utility of additional MSA augmentation. Furthermore, the noise introduced by low-quality MSAs can be amplified in stronger baselines, further impacting performance. 3) Metric Discrepancies: The enhancement effects across different metrics were not entirely consistent. For instance, the trends for pLDDT and RMSD diverged. While pLDDT, as a confidence metric, reflects the model’s confidence in its predictions, RMSD measures the global deviation of the predicted structure from the ground truth. MSA augmentation improved local prediction quality (leading to higher pLDDT), but had limited impact on reducing global structural deviation (resulting in minimal RMSD improvement). Nevertheless, pLDDT alone is sufficient to demonstrate the high quality of the generated MSAs. 4) Noise Issues: EvoDiff and MSAGPT produced MSAs that consistently underperformed compared to the original AF2 MSAs across all modes. This was particularly evident in low-homology protein scenarios, where the generated MSAs failed to accurately capture evolutionary distributions and often included irrelevant sequences. When concatenated with the original AF2 MSAs, these low-quality sequences introduced additional noise, disrupting the model’s inference. This aligns with the hypothesis presented in Section 3.4, where low-quality MSAs were predicted to introduce noise, ultimately impairing folding performance. 5) Few/Zero-shot Consistency: In Few-shot scenarios, EvoDiff and MSAGPT performed relatively better due to the presence of initial MSAs that provided additional homology information to guide the generation process. In contrast, PLAME does not rely on initial MSA searches and directly generates high-quality MSAs through its evolutionary latent space. This enables PLAME to excel in Zero-shot scenarios while maintaining consistent performance in Few-shot cases. This advantage significantly reduces runtime costs while ensuring robust enhancements across all scenarios.
@@ Appendix A Proof of Theorem
Let $P(l,a)$ be the empirical amino–acid distribution for residue $a\!\in\!\mathcal{A}$, and let $Q_{\theta}(l,a)$ denote the model distribution at the residue (i.e. the conditional probability $p_{\theta}(a\mid y_{<l})$ after taking expectation over prefixes). Assign each column a weight $w_{l}\!\in[\,1-\delta,\,1+\delta\,]$ obtained from its conservation score. Then PCE loss directs optimization preferentially toward conserved positions by minimizing a weighted KL divergence and scaling gradient magnitudes in proportion to $w_{l}$.
@@
Re-expressing each column term as $\sum_{a}P\log Q=H\!\bigl{(}P(l,\!\cdot)\bigr{)}+\operatorname{\text{KL}}\!\bigl{(}P(l,\!\cdot)\|Q_{\theta}(l,\!\cdot)\bigr{)}$, we obtain
@@
Based on the fact that $\theta^{\star}_{\mathrm{PCE}}$ minimizes $D_{w}$, denote $\Delta_{w}:=D_{w}(\theta^{\star}_{\mathrm{CE}})-D_{w}(\theta^{\star}_{\mathrm{PCE}})\geq 0$. By applying $(\ast)$ to both optimal parameters and subtracting, we obtain:
@@
Since $\Delta_{w}\geq 0$ and $1-\delta>0$; it is strictly positive whenever $\Delta_{w}>0$, Therefore,
@@
Since $0<\tau<1$, this transformation always increases entropy unless $P$ is already uniform:
@@ Appendix B Training and Sampling Details
Appendix C Extensive Benchmark Results
C.1 Comparison to AI-based MSA Retrieval Methods
We further compare the folding enhancement by MSAs retrieved by AI-based approaches [19]. Without filtering, DHR’s performance is generally higher than AF2 MSA, EvoDiff, and MSAGPT, demonstrating the effectiveness of applying evolutionary embeddings like ESM2.
Zero/Few pLDDT GDT TMscore RMSD LDDT
DHR Zero 63.64 0.51 0.52 12.04 0.55
DHR Few 62.60 0.52 0.57 11.92 0.59
PLAME Zero 66.54 0.53 0.53 11.48 0.57
PLAME Few 66.08 0.54 0.58 12.14 0.60
Table 3: Comparison of folding enhancement to AI-based MSA retrieval method.
C.2 Comparison on Inference Speed and Memory Usage
To further demonstrate PLAME’s efficiency, we calculated the inference time and memory cost of each method. We used ENZYME 1.2.1.50 (EC Number) with length 488 as the test case. The results show that PLAME achieved the fastest speed among all AI-based methods while consuming only 4.5GB of memory. The processing speed is comparable to traditional methods like MMSeq2 and AI-based retrieval methods like DHR. Compared to retrieval-based methods, PLAME does not require downloading or building databases in advance, nor does it need preprocessing steps. This makes it more lightweight and efficient for deployment.
Method Time per MSA (s) GPU Memory (Gb)
PLAME 0.10 4.5
DHR 0.16 + 358.61 (Alignment) 1.9
MMSeq2 0.48 0.0
MSAGPT 62.46 41.6
EvoDiff 478.24 4.0
Table 4: Comparison on inference speed and memory.
C.3 Ablation on HiFiAD Filtering
To demonstrate the effectiveness of HiFiAD, we conducted ablation experiments on filtering methods for other baselines. Among the benchmarks, PLAME maintains SOTA performance, and DHR-HiFiAD and PLAME show comparable performance across multiple metrics. Several metrics of EvoDiff, MSAGPT, and DHR (AI-based MSA retrieval methods) all demonstrate improvement after HiFiAD filtering compared to their original filtering methods, which directly proves the benefit of HiFiAD rules for MSA in AF-series structure prediction. This insight can help future work develop higher-quality MSA sequences.
Zero/Few pLDDT GDT TMscore RMSD LDDT
DHR-HiFiAD Zero 66.01 0.53 0.53 11.48 0.57
DHR-HiFiAD Few 66.08 0.55 0.60 12.14 0.60
EvoDiff-HifiAD Zero 58.24 0.46 0.46 13.74 0.51
EvoDiff-HifiAD Few 60.89 0.49 0.54 12.39 0.56
MSAGPT-HifiAD Zero 60.16 0.48 0.48 12.54 0.53
MSAGPT-HifiAD Few 62.63 0.52 0.57 12.18 0.59
Table 5: Ablation on HiFiAD filtering.
C.4 Ablation on Protein Length
We listed the performance of PLAME and AF2 MSA in different length ranges, and found that PLAME shows overall improvement, with the largest improvement in the 100-300 range, followed by the >>300 and <<100 ranges. We believe this is due to our MSA training data being mainly concentrated in the 100-300 length range.
Length Range pLDDT GDT TMscore RMSD LDDT
AF2 MSA <<100 71.03 0.64 0.52 7.77 0.61
AF2 MSA 100-300 59.50 0.49 0.53 12.46 0.54
AF2 MSA >>300 56.29 0.43 0.51 15.67 0.53
PLAME <<100 74.12 0.63 0.52 7.49 0.61
PLAME 100-300 65.55 0.53 0.58 11.58 0.58
PLAME >>300 58.31 0.45 0.53 16.16 0.54
Table 6: Ablation on protein length.
Appendix D Extensive Case Studies
D.1 Case Study on Successful Designs
To further explore the key pattern of the MSA augmentation, we provide a series of sequence and structure visualization in Appendix G. We select representative cases collected from different datasets and range from different lengths to comprehensive evaluate the samples.